OpenStudy (anonymous):

what is the value of x in sin 3x + sin 2x=0

OpenStudy (xapproachesinfinity):

use sum to product identity \(\large \rm sin(X)+sin(Y)=2sin(\frac{X+Y}{2})cos(\frac{X-Y}{2})\)

OpenStudy (anonymous):

sooooo

OpenStudy (xapproachesinfinity):

you X=3x and your Y=2x so X+Y=5x and X-Y=x then you get \(\Large \rm sin(3x)+cos(2x)=2sin(\frac{5x}{2})cos(\frac{x}{2})\) now we need to solve \(\Large \rm 2sin(5x/2)cos(x/2)=0\) take it from here and solve